# Weight Copy and Low-Rank Adaptation for Few-Shot Distillation of Vision Transformers

### v2

@@ Abstract
Few-shot knowledge distillation recently emerged as a viable approach to harness the knowledge of large-scale pre-trained models, using limited data and computational resources. In this paper, we propose a novel few-shot feature distillation approach for vision transformers. Our approach is based on two key steps. Leveraging the fact that vision transformers have a consistent depth-wise structure, we first copy the weights from intermittent layers of existing pre-trained vision transformers (teachers) into shallower architectures (students), where the intermittence factor controls the complexity of the student transformer with respect to its teacher. Next, we employ an enhanced version of Low-Rank Adaptation (LoRA) to distill knowledge into the student in a few-shot scenario, aiming to recover the information processing carried out by the skipped teacher layers. We present comprehensive experiments with supervised and self-supervised transformers as teachers, on five data sets from various domains, including natural, medical and satellite images. The empirical results confirm the superiority of our approach over competitive baselines. Moreover, the ablation results demonstrate the usefulness of each component of the proposed pipeline.
Keywords:
Knowledge Distillation Low Rank Adaptation Vision Transformers Few-Shot Distillation
@@ 1 Introduction
Vision transformers [4, 19, 32, 45, 54, 69] have revolutionized the computer vision research in the past few years, reaching state-of-the-art performance across a broad range of tasks, such as object recognition [19, 45, 49, 69], object detection [7, 40, 71, 78, 80], image segmentation [10, 11, 20, 23, 73], image translation [33, 56, 61], among many others [32]. Since the accuracy tends to grow as the model gets larger [19, 49], most of the attention has been dedicated to building larger and more powerful models. However, the typically large size and slow inference speed of transformer-based architectures hinders the deployment of such models on environments with limited computational resources, e.g. on edge or mobile devices.
The success of vision transformers lies in the “pre-training then fine-tuning” paradigm, which originates from the natural language processing domain [66, 16]. The multi-head attention layers inside transformers can capture long-range spatial relationships among tokens, but this flexibility has a significant downside: transformers can easily overfit small training sets and experience poor generalization capabilities [37, 44]. Unlike convolutional nets, which benefit from the inductive bias of convolutional filters with small receptive fields [35], transformers requires huge amounts of data in the pre-training stage to avoid overfitting [19], regardless of the chosen supervised [19, 45, 69] or self-supervised [18, 21, 25, 38, 60, 67] training setup. Therefore, to efficiently train large-scale transformer models, powerful and expensive machines are necessary, which are not widely available to researchers. Moreover, large amounts of data are not always available in some domains, e.g. hyperspectral image segmentation [30].
@@
Figure 1: Our feature distillation framework is based on two steps. In the first step, weights from intermittent layers of the teacher transformer are copied directly into the student, where the intermittence factor $r$ coincides with the compression ratio between the teacher and the student transformers. In the second step, enhanced LoRA blocks are integrated into the student network. The enhanced LoRA blocks are trained via feature distillation on unlabeled images. In the illustrated example, the compression ratio is $r=2$.
We emphasize that FSKD has not been extensively explored in the vision domain [24, 39, 41, 42, 51, 55, 59, 76], with even less studies focused on the pre-training stage of vision transformers [42]. To this end, we propose a novel few-shot feature distillation approach for vision transformers based on intermittent Weight Copying and Low-Rank Adaptation (WeCoLoRA), as illustrated in Figure 1. Our approach is divided into two steps. For the first step, we leverage the fact that vision transformers have a consistent depth-wise structure, i.e. the input and output dimensions are compatible across transformer blocks. This allows us to directly copy the weights from intermittent layers of existing pre-trained vision transformers (teachers) into shallower architectures (students). Here, we use the intermittence factor to control the complexity (size) of the student transformer with respect to its teacher. In the second step, we employ an enhanced version of Low-Rank Adaptation (LoRA) [28] to distill knowledge into the student in a few-shot scenario, aiming to recover the information processing carried out by the teacher layers that were skipped in the first step.
We perform the pre-training stage of efficient student transformers via few-shot knowledge distillation on various subsets of ImageNet-1K [15]. Then, we carry out linear probing experiments on five downstream data sets from different image domains, namely ImageNet-1K [15], CIFAR-100 [34], ChestX-ray14 [68], iNaturalist [64] and RESISC45 [12]. We compare WeCoLoRA with state-of-the-art competitors, namely DMAE [3] and DeiT [62], showing that our approach leads to superior results. Furthermore, we conduct an ablation study to demonstrate that each proposed component brings significant performance gains. Additionally, we perform an analysis of the distilled features, which explains why WeCoLoRA produces more robust features than competing methods.
@@
  • We propose a novel few-shot feature distillation approach for vision transformers based on $(i)$ intermittent weight copying and $(ii)$ enhanced low-rank adaption, called WeCoLoRA.
  • We present few-shot and linear probing experiments on five benchmark data sets comprising natural, medical and satellite images, demonstrating the utility of our training pipeline across different domains.
@@ 2 Related Work
Knowledge distillation. Knowledge distillation (KD) [27, 47, 57, 50], a.k.a. teacher-student training, is an efficiency-boosting technique meant to reduce the computational load that comes with large models. It emerged [47] from the union of model compression [27, 2] and learning under privileged information [65]. Our study is only preoccupied with the former task, being aimed at proposing an approach that involves transferring expertise from a heavy teacher model to a lighter student model [6, 63], with the latter finally being able to learn more discriminative features than through a conventional training procedure [14, 29, 13].
KD for vision transformers. KD was originally applied to vision transformers by Touvron et al. [62], who proposed Data-Efficient Image Transformers (DeiT) based on leveraging attention as knowledge and enhancing the student’s capabilities through a learned distillation token. Since then, several other studies used KD on vision transformers [42, 31, 77, 22, 70]. Close to our work, MiniViT [74] uses a parameter reduction technique that involves copying weights (and applying slight perturbations) between the teacher and a smaller student transformer, followed by distillation through self-attention. In comparison, after copying the weights, we employ an enhanced version of LoRA [28] during the information recovery process, which leads to significantly better results in the few-shot scenario. Recently, distilling Masked Autoencoders [25] has gathered some traction [42, 3, 36], with most approaches achieving data efficiency through masking very large percentages (up to 98%) of the input. Our data reduction strategy comes solely from a reduction of the number of unlabeled samples, being more fit for cases where data scarcity is the issue.
@@
Low-Rank Adaptation. LoRA and its variations [28, 81, 72] represent parallel performance improvement methods, enabling the use of powerful large models through lightweight fine-tuning. Consequently, LoRA reduces the inference cost and the requirement to use large data sets. Following other lines of work, the authors of LoRA [28] showed that over-parameterized models reside on a very low internal dimension when making their decisions [52]. The proven hypothesis is that the change in weights during domain adaptation or fine-tuning should also be happening at a low intrinsic rank. To boost training efficiency, we thus combine Low-Rank Adaptation [28] and KD [27] in our study. To the best of our knowledge, there is no prior work that employs weight copying and LoRA as a method for few-shot feature distillation in an unsupervised setting.
@@ 3.1 WeCoLoRA Architecture
Luo et al. [48] proposed to apply LoRA only to the projection matrices corresponding to the query and value tokens $\boldsymbol{Q}$ and $\boldsymbol{V}$, respectively. However, in order to fully replicate the functionality of the teacher, we conjecture that it is better to apply LoRA to each component of the transformer block. Therefore, for each head $h_{i}$, $\forall i\in\{1,2,...,H\}$ of the multi-head self-attention (MSA) layer, the outputs of the query $\boldsymbol{Q}_{i}$, value $\boldsymbol{V}_{\!i}$ and key $\boldsymbol{K}_{i}$ applied to the input $x$ become:
@@
The formula to compute the output of the head $h_{i}$ remains unaltered, being equal to $h_{i}=Attention(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{\!i})$, $\forall i\in\{1,2,...,H\}$, where:
$Attention(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\mbox{softmax}\left(\frac{\boldsymbol{Q}\cdot\boldsymbol{K}^{\top}}{\sqrt{d_{k}}}\right)\cdot\boldsymbol{V},$ (2)
@@
$\begin{split}MultiHead(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})&=\boldsymbol{W}_{\!O}\cdot[h_{1},h_{2},..,h_{H}]+\boldsymbol{B}_{O}\cdot\boldsymbol{A}_{O}\cdot[h_{1},h_{2},..,h_{H}],\end{split}$ (3)
where FFL is the output of the first layer in the FFN model, $\boldsymbol{W}_{\!f_{1}}$ and $\boldsymbol{W}_{\!f_{2}}$ are the parameters of the feed-forward network, $\sigma$ is the activation function, and $\boldsymbol{A}_{*}$ and $\boldsymbol{B}_{*}$ are the matrices learned by LoRA, where $*\in\{f_{1},f_{2}\}$. The biases in the FFN network were omitted to enhance the clarity of the presentation. By introducing LoRA in every component of the transformer block, we obtain an enhanced version of LoRA-based transformer blocks.
@@ 4 Experiments
We use ImageNet-1K [15] to perform knowledge distillation in various few-shot setups, considering subsets ranging between $1\%$ and $10\%$ of the number of samples in the original training set. We test the models via linear probing, using the official ImageNet-1K evaluation set. We also perform linear probing experiments on five downstream data sets: ImageNet-1K [15], CIFAR-100 [34], ChestX-ray14 [68], iNaturalist [64] and RESISC45 [12].
@@ 4.1 Data Sets
CIFAR-100. CIFAR-100 [34] is a balanced data set of 60,000 images representing various visual entities, grouped in 100 classes. The categories are arranged in a hierarchical structure, under 20 superclasses. The official test set contains 10,000 pictures. Each color image has $32\times 32$ pixels.
@@
iNaturalist. iNaturalist [64] is a heavily imbalanced data set, proposing the fine-grained classification of 1,010 plant and animal species. Comprising a total size of 268,243 images with varying resolutions, its 2019 version offers an evaluation set of 3030 images.
@@ 4.2 Implementation Details
We implement our method in Pytorch. We create the enhanced version of LoRA starting from the official code of LoRA [^1] https://github.com/microsoft/LoRA and extending it, in order to be applied to each component of the transformer layer. To demonstrate the generalization of our method to different teachers, we perform experiments with a supervised and a self-supervised teacher, respectively. As the supervised teacher backbone, we employ the ViT-B model [19] trained on ImageNet, available in the timm [^2] https://pypi.org/project/timm library. The selected self-supervised model is also based on ViT-B, but pre-trained using the Masked Autoencoder framework [25]. Since the goal is to create a model which is faster and lighter, we refrain from using larger teacher models, which would be inefficient for our purpose.
To show that our method generalizes to different depth reduction factors, we perform experiments with $r\!=\!2$ and $r\!=\!3$. Since the number of layers $L$ in the teacher architecture (ViT-B) is $12$, the number of layers in the student architecture becomes $M\!=\!6$ (when $r\!=\!2$) or $M\!=\!4$ (when $r\!=\!3$), respectively. We select the rank of the low-rank matrices from $k\in\{128,256,512\}$, which controls the number of parameters used to recover the skipped layers from the teacher.
We perform the knowledge distillation procedure for only $10$ epochs. We set the learning rate to $10^{-3}$ when the compression factor $r$ equals $2$, and to $10^{-4}$ when $r$ equals $3$. The batch size is set to $128$, with a gradient accumulation factor of $8$. We perform weak data augmentation, similar to He et al. [25].
@@ 4.3 Results
$\alpha$ Weight Copy LoRA Enhanced LoRA ImageNet iNaturalist CIFAR-100
@@
Table 2: Accuracy rates (in percentages) obtained on ImageNet-1K [15], iNaturalist19 [64] and CIFAR-100 [34] by two state-of-the-art knowledge distillation methods, DMAE [3] and DeiT [62], in comparison with WeCo+KD and WeCoLoRA. The results are obtained using a compression factor of $r=2$, employing the self-supervised teacher ViT-B [25]. The parameter $\alpha$ represents the percentage of the original ImageNet-1K training set [15] used during knowledge distillation. The best accuracy on each downstream data set and each value of $\alpha$ is highlighted in bold.
$\alpha$ Method Venue ImageNet iNaturalist CIFAR-100
$\alpha=1$ DMAE [3] CVPR 2023 $15.5$ $8.5$ $22.4$
DeiT [62] ICML 2021 $16.5$ $12.9$ $28.8$
WeCo+KD (ours, ablated) - $30.2$ $19.2$ $32.1$
WeCoLoRA (ours) - $\mathbf{36.8}$ $\mathbf{22.7}$ $\mathbf{33.3}$
$\alpha=10$ DMAE [3] CVPR 2023 $30.0$ $17.1$ $31.7$
DeiT [62] ICML 2021 $39.9$ 29.2 $48.1$
WeCo+KD (ours, ablated) - $53.7$ $31.2$ $48.2$
WeCoLoRA (ours) - $\mathbf{56.0}$ $\mathbf{34.5}$ $\mathbf{49.8}$
- ViT-B [25] (teacher) CVPR 2022 $66.1$ $41.6$ $55.1$
Comparison with state-of-the-art KD techniques. We compare WeCoLoRA and WeCo+KD with two state-of-the-art knowledge distillation techniques, namely DMAE [3] and DeiT [62], in Table 5. We also report the performance obtained by the self-supervised teacher ViT-B [25] on each downstream task, as an upper bound for the few-shot methods. Interestingly, neither state-of-the-art method [3, 62] outperforms the ablated version of WeCoLoRA, namely WeCo+KD. This indicates that current methods, such as DMAE and DeiT, are not able to handle the limited amount of data that is typical to the few-shot setting. WeCoLoRA surpasses the two state-of-the-art KD methods [3, 62] by significant margins, e.g. the differences are between $16\%$ and $26\%$ on ImageNet-1K. Moreover, on the CIFAR-100 data set, our method obtains an accuracy that is only $5.3\%$ below the teacher, while using half the number of layers and only using $10\%$ of the training data during distillation. Since WeCo+KD, the ablated version of our method, surpasses both DMAE [3] and DeiT [62], we choose it as a strong baseline for the subsequent experiments.
@@
We present results for $10\%$ of the distillation data in Table 4. Once again, WeCoLoRA outperforms the baseline in most settings. However, the gains in favor of our method when using $10\%$ of the data (Table 4) are generally lower than distilling on $1\%$ of the data (Table 3). We consider that $10\%$ of ImageNet-1K, i.e. about 100K training images, is at the upper end of the scenarios that can still be regarded as few-shot. The higher number of trainable parameters in WeCo+KD, combined with the large training set, allow it to recover the gap with respect to WeCoLoRA. To further demonstrate the superiority of WeCoLoRA in other few-shot scenarios, we present additional results when using $2\%$ and $5\%$ of the data during distillation, in the supplementary.
@@
Latent space visualizations. In Figure 2, we illustrate t-SNE visualizations of the embeddings learned with the proposed method (WeCoLoRA) versus its ablated version denoted as WeCo+KD. The embeddings are obtained on the RESISC45 [12] test set. We illustrate the embeddings of the student obtained through distilling the supervised ViT-B [19] teacher, with a compression ratio set to $r=2$. We notice that our WeCoLoRA is able to learn a discriminative feature distribution during distillation, disentangling the out-of-domain (downstream) data samples from RESISC45 into clearly delineated clusters, even though there was no supervision involved, other than the teacher features. However, WeCo+KD, which is based on complete weight adaptation during distillation, creates clusters which are spread across the feature space. In summary, the t-SNE visualizations emphasize that WeCoLoRA is an effective few-shot knowledge distillation approach, explaining the superior results on downstream linear probing tasks through the robustness of the learned latent space.
@@
Attention visualizations. To visualize the attention, we employ Attention Rollout [1], a technique that enables the visualization of the attention flow throughout ViT. We fuse the attention heads by taking the maximum response (top $10\%$) and discard pixels with lower values. In Figure 3, we present three test images that are randomly taken from ImageNet [15], which correspond to “Daisy”, “Persian Cat” and “Goldfish” classes, respectively. The attention is extracted from student models trained with WeCo+KD and WeCoLoRA. We observe that WeCoLoRA makes the model focus more on discriminative features such as petals, fur, and fish scales. This helps the model based on WeCoLoRA in taking informed decisions.
@@
5 Limitations
The main limitation of our method is its applicability to vision transformers and architectures that use multiple consecutive blocks with the same configuration, e.g. ResNets [26]. This restriction is imposed by our weight copying mechanism. Our ablation results indicate that the weight copying step is very useful in the few-shot distillation scenario, as it significantly boosts performance (see Table 1). Simply removing the weight copying step is not a viable option, since the performance would drastically degrade. To make our framework applicable to any architecture, the weight copying mechanism could be enhanced with adaptor blocks, which would be able to reshape the copied weights to the appropriate size. However, the adaptor blocks need to be tailored for each specific pair of teacher and student models. This will increase the complexity of the hyperparameter tuning stage, which, in the current form, is quite straightforward, i.e. aside from typical hyperparameters, such as the learning rate and the mini-batch size, WeCoLoRA only adds the compression ratio $r$ and the rank of the low-rank matrices $k$ as extra hyperparameters.

### v3

@@ Abstract
Few-shot knowledge distillation recently emerged as a viable approach to harness the knowledge of large-scale pre-trained models, using limited data and computational resources. In this paper, we propose a novel few-shot feature distillation approach for vision transformers. Our approach is based on two key steps. Leveraging the fact that vision transformers have a consistent depth-wise structure, we first copy the weights from intermittent layers of existing pre-trained vision transformers (teachers) into shallower architectures (students), where the intermittence factor controls the complexity of the student transformer with respect to its teacher. Next, we employ an enhanced version of Low-Rank Adaptation (LoRA) to distill knowledge into the student in a few-shot scenario, aiming to recover the information processing carried out by the skipped teacher layers. We present comprehensive experiments with supervised and self-supervised transformers as teachers, on six data sets from various domains (natural, medical and satellite images) and tasks (classification and segmentation). The empirical results confirm the superiority of our approach over state-of-the-art competitors. Moreover, the ablation results demonstrate the usefulness of each component of the proposed pipeline. We release our code at https://github.com/dianagrigore/WeCoLoRA.
@@ 1 Introduction
Figure 1: Our unsupervised feature distillation framework is based on two steps. In the first step, weights from intermittent layers of the teacher transformer are copied directly into the student, where the intermittence factor $r$ coincides with the compression ratio between the teacher and the student transformers. In the second step, enhanced LoRA blocks are integrated into the student network. The enhanced LoRA blocks are trained via feature distillation on unlabeled images. In the illustrated example, the compression ratio is $r=2$. Best viewed in color.
Vision transformers [5, 20, 33, 47, 56, 74] have revolutionized computer vision research in the past few years, reaching state-of-the-art performance across a broad range of tasks, such as object recognition [20, 47, 51, 74], object detection [8, 41, 76, 83, 85], image segmentation [11, 12, 21, 24, 78], image translation [34, 58, 66], among many others [33]. Since the accuracy tends to grow as the model gets larger [20, 51], most of the attention has been dedicated to building larger and more powerful models. However, the typically large size and slow inference speed of transformer-based architectures hinders the deployment of such models on environments with limited computational resources, e.g. on edge or mobile devices.
The success of vision transformers lies in the “pre-training then fine-tuning” paradigm, which originates from the natural language processing domain [71, 17]. The multi-head attention layers inside transformers can capture long-range spatial relationships among tokens, but this flexibility has a significant downside: transformers can easily overfit small training sets and experience poor generalization capabilities [38, 46]. Unlike convolutional nets, which benefit from the inductive bias of convolutional filters with small receptive fields [36], transformers requires huge amounts of data in the pre-training stage to avoid overfitting [20], regardless of the chosen supervised [20, 47, 74] or self-supervised [19, 22, 26, 39, 65, 72] training setup. Therefore, to efficiently train large-scale transformer models, powerful and expensive computers are necessary, which are not widely available to researchers. Moreover, large amounts of data are not always available in some domains, e.g. hyperspectral image segmentation [31].
@@
We emphasize that FSKD has not been extensively explored in the vision domain [25, 40, 42, 43, 53, 57, 64, 81], with even less studies focused on the pre-training stage of vision transformers [43]. To this end, we propose a novel few-shot feature distillation approach for vision transformers based on intermittent Weight Copying and Low-Rank Adaptation (WeCoLoRA), as illustrated in Figure 1. Our approach is divided into two steps. For the first step, we leverage the fact that vision transformers have a consistent depth-wise structure, i.e. the input and output dimensions are compatible across transformer blocks. This allows us to directly copy the weights from intermittent layers of existing pre-trained vision transformers (teachers) into shallower architectures (students). Here, we use the intermittence factor to control the complexity (size) of the student transformer with respect to its teacher. In the second step, we employ an enhanced version of Low-Rank Adaptation (LoRA) [29] to distill knowledge into the student in a few-shot scenario, aiming to recover the information processing carried out by the teacher layers that were skipped in the first step. We hereby note that standard LoRA is usually employed over queries and values, which is insufficient to fully replicate the behavior of the skipped blocks. To mitigate this issue and better complement the weight copying mechanism, we apply enhanced LoRA, a variant of LoRA that approximates all layers in the skipped transformer block.
We perform the pre-training stage of efficient student transformers via few-shot knowledge distillation on various subsets of ImageNet-1K [16]. Then, we carry out linear probing experiments on five downstream data sets from different image domains, namely ImageNet-1K [16], CIFAR-100 [35], ChestX-ray14 [73], iNaturalist [69] and RESISC45 [13]. We compare WeCoLoRA with state-of-the-art competitors, namely DeiT [67], DMAE [4], MiniViT [79] and DoRA [45], showing that our approach leads to superior results. Furthermore, we conduct an ablation study to demonstrate that each proposed component brings significant performance gains. Additionally, we perform an analysis of the distilled features, which explains why WeCoLoRA produces more robust features than competing methods. In the supplementary, we demonstrate the versatility of WeCoLoRA by applying it on a convolutional network for image segmentation on the MSLesSeg [60] data set.
@@
  • We propose a novel unsupervised few-shot feature distillation approach for vision transformers based on $(i)$ intermittent weight copying and $(ii)$ enhanced low-rank adaption, called WeCoLoRA.
  • We present few-shot and linear probing experiments on six benchmark data sets comprising natural, medical and satellite images, demonstrating the utility of our training pipeline across different domains and tasks.
@@ 2 Related Work
Knowledge distillation. Knowledge distillation (KD) [28, 49, 59, 52], a.k.a. teacher-student training, is an efficiency boosting technique meant to reduce the computational load that comes with large models. It emerged [49] from the union of model compression [28, 3] and learning under privileged information [70]. Our study is only preoccupied with the former task, being aimed at proposing an approach that involves transferring expertise from a heavy teacher model to a lighter student model [7, 68], with the latter finally being able to learn more discriminative features than through a conventional training procedure [15, 30, 14].
KD for vision transformers. KD was originally applied to vision transformers by Touvron et al. [67], who proposed Data-Efficient Image Transformers (DeiT) based on leveraging attention as knowledge and enhancing the student’s capabilities through a learned distillation token. Since then, several other studies used KD on vision transformers [43, 32, 82, 23, 75]. Close to our work, MiniViT [79] uses a parameter reduction technique that involves copying weights (and applying slight perturbations) between the teacher and a smaller student transformer, followed by distillation through self-attention. In comparison, after copying the weights, we employ an enhanced version of LoRA [29] during the information recovery process, which leads to significantly better results in the few-shot scenario. Recently, distilling Masked Autoencoders [26] has gathered some traction [43, 4, 37], with most approaches achieving data efficiency through masking very large percentages (up to 98%) of the input. Our data reduction strategy comes solely from a reduction of the number of unlabeled samples, being more fit for cases where data scarcity is an issue.
@@
Low-Rank Adaptation. LoRA and its variations [29, 45, 86, 77] represent parallel performance improvement methods, enabling the use of powerful large models through lightweight fine-tuning. Consequently, LoRA reduces the inference cost and the requirement to use large data sets. Following other lines of work, the authors of LoRA [29] showed that over-parameterized models rely on a very low internal dimension when making their decisions [54]. The proven hypothesis is that the change in weights during domain adaptation or fine-tuning should also be happening at a low intrinsic rank. To boost training efficiency, we thus combine Low-Rank Adaptation [29] and KD [28] in our study. To the best of our knowledge, there is no prior work that employs weight copying and LoRA as a method for few-shot feature distillation in an unsupervised setting.
@@ 3.1 WeCoLoRA Architecture
Luo et al. [50] proposed to apply LoRA only to the projection matrices corresponding to the query and value tokens $\boldsymbol{Q}$ and $\boldsymbol{V}$, respectively. However, in our setup, after applying the intermittent weight copying mechanism, we need to replicate the functionality of the skipped transformer blocks. Since LoRA only adapts the projection matrices of query and value tokens, it does not provide the means to fully replicate the skipped blocks. In order to mitigate this issue, we conjecture that it is better to apply LoRA to each component of the transformer block. Therefore, for each head $h_{i}$, $\forall i\in\{1,2,...,H\}$ of the multi-head self-attention (MSA) layer, the outputs of the query $\boldsymbol{Q}_{i}$, value $\boldsymbol{V}_{\!i}$ and key $\boldsymbol{K}_{i}$ applied to the input $x$ become:
@@
The formula to compute the output of the head $h_{i}$ remains unaltered, being equal to $h_{i}=\mbox{Attention}(\boldsymbol{Q}_{i},\boldsymbol{K}_{i},\boldsymbol{V}_{\!i})$, $\forall i\in\{1,2,...,H\}$, where:
$\mbox{Attention}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\mbox{softmax}\!\left(\frac{\boldsymbol{Q}\cdot\boldsymbol{K}^{\top}}{\sqrt{d_{k}}}\right)\cdot\boldsymbol{V},$ (2)
@@
$\begin{split}\mbox{MultiHead}(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})&=\boldsymbol{W}_{\!O}\cdot[h_{1},h_{2},..,h_{H}]+\\ &+\boldsymbol{B}_{O}\cdot\boldsymbol{A}_{O}\cdot[h_{1},h_{2},..,h_{H}],\end{split}$ (3)
where FFL is the output of the first layer in the FFN model, $\boldsymbol{W}_{\!f_{1}}$ and $\boldsymbol{W}_{\!f_{2}}$ are the parameters of the feed-forward network, $\sigma$ is the activation function, and $\boldsymbol{A}_{*}$ and $\boldsymbol{B}_{*}$ are the matrices learned by LoRA, where $*\in\{f_{1},f_{2}\}$. The biases in the FFN network were omitted to enhance the clarity of the presentation. By introducing LoRA in every component of the transformer block, we obtain an enhanced version of LoRA-based transformer blocks. Enhanced LoRA is able to better complement the weight copying mechanism by enabling a more faithful replication of the skipped teacher blocks. This statement is confirmed by our ablation study.
@@ 4 Experiments
We use ImageNet-1K [16] to perform knowledge distillation in various few-shot setups, considering subsets ranging between $0.25\%$ and $10\%$ of the number of samples in the original training set. We test the models via linear probing, using the official ImageNet-1K evaluation set. We also perform linear probing experiments on five downstream data sets: ImageNet-1K [16], CIFAR-100 [35], ChestX-ray14 [73], iNaturalist [69] and RESISC45 [13].
@@ 4.1 Data Sets
CIFAR-100. CIFAR-100 [35] is a balanced data set of 60,000 images representing various visual entities, grouped into 100 classes. The categories are arranged in a hierarchical structure, under 20 superclasses. The official test set contains 10,000 pictures. Each image has $32\times 32$ pixels.
@@
iNaturalist. iNaturalist [69] is a heavily imbalanced data set, proposing the fine-grained classification of 1,010 plant and animal species. Comprising a total size of 268,243 images with varying resolutions, its 2019 version offers an evaluation set of 3,030 images.
@@ 4.2 Implementation Details
We implement our method in Pytorch. We create the enhanced version of LoRA starting from the official code of LoRA [^1] https://github.com/microsoft/LoRA and extending it, in order to be applied to each component of the transformer layer. To demonstrate the generalization of our method to different teachers, we perform experiments with a supervised and a self-supervised teacher, respectively. As the supervised teacher backbone, we employ the ViT-B model [20] trained on ImageNet, available in the timm [^2] https://pypi.org/project/timm library. The selected self-supervised model is also based on ViT-B, but pre-trained using the Masked Autoencoder framework [26]. Since the goal is to create a model which is faster and lighter, we generally refrain from using larger teacher models, which would be inefficient for our purpose.
To show that our method generalizes to different depth reduction factors, we perform experiments with $r\!=\!2$, $r\!=\!3$ and $r\!=\!4$. Since the number of layers $L$ in the teacher architecture (ViT-B) is $12$, the number of layers in the student architecture becomes $M\!=\!6$ (when $r\!=\!2$), $M\!=\!4$ (when $r\!=\!3$), or $M\!=\!3$ (when $r\!=\!4$), respectively. We select the rank of the low-rank matrices from $k\in\{128,256,512\}$, which controls the number of parameters used to recover the skipped layers from the teacher.
We perform the knowledge distillation procedure for only $10$ epochs. We set the learning rate to $10^{-3}$ when the compression factor $r$ equals $2$, and to $10^{-4}$ when $r$ is larger. The batch size is set to $128$, with a gradient accumulation factor of $8$. We perform weak data augmentation, similar to He et al. [26].
@@ 4.3 Results
$\alpha$
Weight Copy
    LoRA
Enhanced LoRA
ImageNet
iNaturalist
CIFAR-100
@@
$\alpha$ Method Venue Self-supervised teacher Supervised teacher
ImageNet iNaturalist CIFAR-100 ImageNet iNaturalist CIFAR-100
$\alpha=1$ DeiT [67] ICML 2021 $16.5$ $12.9$ $28.8$ $25.5$ $20.3$ $39.0$
MiniViT [79] CVPR 2022 - - - $26.7$ $21.0$ $37.3$
DMAE [4] CVPR 2023 $15.5$ $8.5$ $22.4$ $15.4$ $7.9$ $22.2$
DoRA [45] ICML 2024 - - - $29.0$ $20.9$ $25.6$
WeCo+KD (ours, ablated) - $30.2$ $19.2$ $32.1$ $35.4$ $25.4$ $34.8$
WeCoLoRA (ours) - $\mathbf{36.8}$ $\mathbf{22.7}$ $\mathbf{33.3}$ $\mathbf{60.5}$ $\mathbf{43.4}$ $\mathbf{60.6}$
$\alpha=10$ DeiT [67] ICML 2021 $39.9$ $29.2$ $48.1$ $57.6$ $37.8$ $61.4$
MiniViT [79] CVPR 2022 - - - $64.0$ $40.8$ $64.2$
DMAE [4] CVPR 2023 $30.0$ $17.1$ $31.7$ $33.0$ $21.0$ $35.6$
DoRA [45] ICML 2024 - - - $58.7$ $45.3$ $58.7$
WeCo+KD (ours, ablated) - $53.7$ $31.2$ $48.2$ $68.3$ $45.7$ $64.4$
WeCoLoRA (ours) - $\mathbf{56.0}$ $\mathbf{34.5}$ $\mathbf{49.8}$ $\mathbf{69.2}$ $\mathbf{49.5}$ $\mathbf{68.3}$
$\alpha=100$ ViT-B (teacher) - $66.1$ $41.6$ $55.1$ $84.2$ $52.5$ $89.1$
Table 2: Accuracy rates (in percentages) obtained on ImageNet-1K [16], iNaturalist19 [69] and CIFAR-100 [35] by four state-of-the-art methods, DeiT [67], MiniViT [79], DMAE [4] and DoRA [45], in comparison with WeCo+KD and WeCoLoRA. The results are obtained using a compression factor of $r=2$, while alternating between the self-supervised ViT-B teacher [26] and the supervised ViT-B teacher [20]. The parameter $\alpha$ represents the percentage of the original ImageNet-1K training set [16] used during knowledge distillation. The best accuracy on each downstream data set and each value of $\alpha$ is highlighted in bold. The teachers are based on linear probing.
@@
Comparison with state-of-the-art methods. In Table 2, we compare WeCoLoRA and WeCo+KD with four state-of-the-art techniques: DeiT [67], MiniViT [79], DMAE [4] and DoRA [45]. We also report the performance obtained by the self-supervised [26] and supervised [20] teachers (both based on ViT-B) on each downstream task, as an upper bound for the few-shot methods. Note that MiniViT and DoRA are not directly applicable in conjunction with self-supervised teachers.
Interestingly, the ablated version of WeCoLoRA, namely WeCo+KD, generally outperforms all state-of-the-art methods [4, 45, 67, 79]. This indicates that current methods are not able to handle the limited amount of data that is typical to the few-shot setting. WeCoLoRA surpasses all four state-of-the-art KD methods [4, 45, 67, 79] by significant margins, e.g. the differences are between $5\%$ and $45\%$ on ImageNet-1K. Moreover, on the iNauralist data set, our method obtains an accuracy that is only $3\%$ below the supervised teacher, while using half the number of layers and only using $10\%$ of the training data during distillation. Since WeCo+KD, the ablated version of our method, generally surpasses the state-of-the-art methods, we choose it as a strong baseline for the subsequent experiments.
@@
We present results for $10\%$ of the distillation data in Table 4. Once again, WeCoLoRA outperforms the baseline in most settings. However, the gains in favor of our method when using $10\%$ of the data (Table 4) are generally lower than distilling on $1\%$ of the data (Table 3). We consider that $10\%$ of ImageNet-1K, i.e. about 100K training images, is at the upper end of the scenarios that can still be regarded as few-shot. The higher number of trainable parameters in WeCo+KD, combined with the large training set, allow it to recover the gap with respect to WeCoLoRA. To further demonstrate the superiority of WeCoLoRA in other few-shot scenarios, we present additional results when using $0.25\%$, $2\%$ and $5\%$ of the data during distillation, in the supplementary.
@@
Latent space visualizations. In Figure 2, we illustrate t-SNE visualizations of the embeddings learned with the proposed method (WeCoLoRA) versus its ablated version denoted as WeCo+KD. The embeddings are obtained on the RESISC45 [13] test set. We illustrate the embeddings of the student obtained through distilling the supervised ViT-B [20] teacher, with a compression ratio of $r=2$. We notice that our WeCoLoRA is able to learn a discriminative feature distribution during distillation, disentangling the out-of-domain (downstream) data samples from RESISC45 into clearly delineated clusters, even though there was no supervision involved, other than the teacher features. However, WeCo+KD, which is based on complete weight adaptation during distillation, creates clusters which are spread across the feature space. In summary, the t-SNE visualizations emphasize that WeCoLoRA is an effective few-shot knowledge distillation approach, explaining the superior results on downstream linear probing tasks through the robustness of the learned latent space.
@@
Attention visualization. To visualize the attention, we employ Attention Rollout [1], a technique that enables the visualization of the attention flow throughout ViT. We fuse the attention heads by taking the maximum response (top $10\%$) and discard pixels with lower values. In Figure 3, we present three test images that are randomly taken from ImageNet [16], which correspond to “Daisy”, “Persian Cat” and “Goldfish” classes, respectively. The attention is extracted from student models trained with WeCo+KD and WeCoLoRA. We observe that WeCoLoRA makes the model focus more on discriminative features such as petals, fur, and fish scales. This helps the model based on WeCoLoRA in taking informed decisions.